OpenStudy (anonymous):

What is the ∫1/x2+x+1 dx ?

OpenStudy (anonymous):

You completed the square perfectly fine. Now the integral looks like this: \(\large \int \frac{1}{(x+\frac{1}{2})^2+(\frac{\sqrt 3}{2})^2} dx\)

OpenStudy (anonymous):

Why do you put 3/4 as the (3√2)^2 ? \[ \int\limits_{}^{} \frac{\frac{ 4 }{ 3 }\ 1 }{ (\frac{ 4 }{ 3 }((x+\frac{ 1 }{ 2 })^{2} +\frac34)} \] \[\frac{ 4 }{ 3 }\ \int\limits_{}^{} \frac{ 1 }{ (\frac{ 4 }{ 3 }(x+\frac{ 1 }{ 2 })^{2} +1)} \] And what is next ?

OpenStudy (anonymous):

This 3/4 has been written as (sqrt(3/4))^2 because we know a formula which says...  Ofcourse with a constant :-)